OpenStudy (anonymous):

Help!! The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2).

OpenStudy (anonymous):

I think it's y=-1/2x+3/2

OpenStudy (dan815):

you hae to take the negative recripcoal so if you are given a slope 2 or 2/1 then u need -1/2 as the perpendicular slope or if u are given -2 or -2/1 then -(-1/2) = 1/2 slope

OpenStudy (dan815):

from that you know the slope has to be 1/2 y=1/2 x + b can you solve for b , given the point (-1,-2)

OpenStudy (anonymous):

How would I plug those in?

OpenStudy (anonymous):

Or how do I solve it?

OpenStudy (dan815):

a point is giving you a y value for some x value -1,-2 means at x=-1 y=-2

OpenStudy (dan815):

y=1/2x+b sub in the point (-1,-2) -2=1/2(-1)+b now solve for b

OpenStudy (anonymous):

So it's y=-1/2x-3/2

OpenStudy (dan815):

i toold you why that cant be xD

OpenStudy (anonymous):

Oh sorry y=1/2x-3/2?

OpenStudy (dan815):

yep

OpenStudy (anonymous):

Thank you so much, can you check another one for me?

OpenStudy (dan815):

ok sure

OpenStudy (anonymous):

Line QR contains (2, 8) and (3, 10) Line ST contains points (0, 6) and (−2, 2). Lines QR and ST are I think the answer is "Parallel, because the slopes are the same"

OpenStudy (dan815):

can u show me how u are finding the slope for these 2 lines

OpenStudy (anonymous):

For QR I have y2-y1/x2-x1. 10-8/3-2 = 2/1 = 2 For ST I have 2-6/(-2)-0= -4/-2= 2

OpenStudy (dan815):

good :)

OpenStudy (dan815):

you are right

OpenStudy (anonymous):

And they are parallel not perpendicular right?

OpenStudy (dan815):

ya same slope

OpenStudy (dan815):

y=mx+b if they have the same slope, then the only difference will be the y intercept so it will be the same line shifted up or down

OpenStudy (anonymous):

I have an answer choice that says the same thing except perpendicular

OpenStudy (dan815):



OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (dan815):



OpenStudy (dan815):

oops typo

OpenStudy (dan815):

oh you are welcome :)